OpenStudy (anonymous):

How do you find the inverse of the equation y= x/1-x

OpenStudy (anonymous):

\[x=\frac{y}{y-1}\] solve for \(y\) first step is \[x(y-1)=y\] then multiply out to get \[xy-x=y\] then \[xy-y=x\] and \[(x-1)y=x\] one more step

OpenStudy (anonymous):

in fact you get \[y=\frac{x}{x-1}\] so it is its own inverse

OpenStudy (anonymous):

i actually did it out and i think it's y=x/(x+1)